OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (hlares):

To solve this system of equations, you will want to first eliminate one of the variables by adding the two equations together, thus: x-2y=14 Multiply by 3 x+3y=9 Multiply by 2 --------- 3x-6y=42 Add 2x+6y=18 --------- 5x=60 x=12 Now that you have x, you just need to substitute it back into the first equation and solve for y: (12)-2y=14 -2y=2 y=-1 And there you go! Your answer is (12, -1).

OpenStudy (anonymous):

(1, 12) (−1, −12) y=-1